# Neural Post-Einsteinian Test of General Relativity with the Third Gravitational-Wave Transient Catalog

### v1

@@ Acknowledgements.
Acknowledgments. Y.X. and N.Y. acknowledge support from the Simons Foundation through Award No. 896696, the NSF through Grant No. PHY-2207650 and NASA through Grant No. 80NSSC22K0806. Y.X. also acknowledges support from the Illinois Center for Advanced Studies of the Universe (ICASU)/Center for AstroPhysical Surveys (CAPS) Graduate Fellowship. G. N. acknowledges NSF support from AST-2206195, and a CAREER grant, supported in-part by funding from Charles Simonyi, NSF AST 2421845 and support from the Simons Foundation as part of the NSF-Simons SkAI Institute. This work made use of the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA), and is supported by funds from the University of Illinois Urbana-Champaign (UIUC).
@@
With properly chosen priors, Bayes’ theorem can be applied to the hierarchical likelihood $\mathcal{L}_{\rm h}$ to obtain the posterior distribution for the hyperparameters, $p_{\rm h}(\mu,\sigma,\bar{\varphi}|\{s\})$, based on which the GR quantile can be introduced as [85],
@@
Table 1: Events selected for our analysis and the list of detectors operated during each event. For the latter, the abbreviations “H,” “L,” and “V” correspond to the Hanford detector, the Livingston detector, and the Virgo detector, respectively. As pointed out later, three events fail the consistency check when comparing the npE posteriors from different detectors, suggesting certain noise artifacts that may be significantly affecting the inference process. These events are removed from the results presented in the main text. In addition, the strain data for GW190412 is known to have significant contributions from higher harmonics, and so we take special care when modeling that signal.
By default, we choose IMRPhenomPv2 as the base GR waveform $\tilde{h}_{\rm GR}$. However, in the special case of GW190412, we choose IMRPhenomXPHM with an additional $(3,3)$ mode added on top of the dominant $(2,2)$ mode, as mentioned in the caption of Table 1. Both GR waveforms are parametrized by
@@
Figure 3: Posterior 90% credible contours from individual-event npE analysis assuming a uniform prior over $\zeta_{1}$ and $\zeta_{2}$ within the unit circle. Each panel presents the analysis result of an event in Table 1, as suggested by the panel title, and apart from the posterior generated using data of the entire network (black), there are also posteriors generated using data of individual detectors including the Hanford detector (blue), the Livingston detector (orange), and the Virgo detector (green). See Table 1 for the list of detectors used in each event. The ticks and labels of the panel axes have been omitted for simplicity. For all panels, the $x$-axis represents $\zeta_{1}\in[-1,1]$ and the $y$-axis represents $\zeta_{2}\in[-1,1]$. Each panel is also gridded by gray dotted lines, which are the same PN lines shown in Fig. 2 in the main text.
@@
For the rest of the events, the network posteriors are mostly consistent with GR by enclosing $\vec{\zeta}=0$ inside the black 90% credible contours. Exceptions are GW191129_134029 and GW190924_021846, but their network 90% credible contours are not far from the GR point. There are also a few events such as GW150914, where the enclosure is only critically fulfilled. However, these cases should be seen as from the tail of the GW population, and do not necessarily indicate a break of GR at the catalog level. This argument has been further strengthened by the hierarchical npE test result, from which we have presented a $Q_{\rm GR}=34\%<90\%$ in the main text.
@@
Figure 4: Hierarchical posterior of the population hyperparameters. The blue contours enclose the 50% and 90% credible regions, respectively.

### v2

@@ Acknowledgements.
Acknowledgments. Y.X. and N.Y. acknowledge support from the Simons Foundation through Award No. 896696, the Simons Foundation International through Award No. SFI-MPS-BH-00012593-01, the NSF through Grants No. PHY-2207650 and No. PHY-25-12423, and NASA through Grant No. 80NSSC22K0806. Y.X. also acknowledges support from the Illinois Center for Advanced Studies of the Universe (ICASU)/Center for AstroPhysical Surveys (CAPS) Graduate Fellowship and support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [funding reference number 513671]. G. N. acknowledges NSF support from AST-2206195, and a CAREER grant, supported in part by funding from Charles Simonyi, NSF AST 2421845 and support from the Simons Foundation as part of the NSF-Simons SkAI Institute. This work made use of the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA), and is supported by funds from the University of Illinois Urbana-Champaign (UIUC).
Data availability. The data that support the findings of this article are openly available [75, 76, 104, 105, 106].
@@
With priors specified, Bayes’ theorem can be applied to the hierarchical likelihood $\mathcal{L}_{\rm h}$ to obtain the posterior distribution for the hyperparameters, $p_{\rm h}(\mu,\sigma,\bar{\varphi}|\{s\})$, based on which the GR quantile can be introduced as [101],
@@
GW241011_233834 HV Higher harmonics [60]
GW250114_082203 HL
Table 1: Events selected for our analysis and the list of detectors operated during each event. For the latter, the abbreviations “H,” “L,” and “V” correspond to the Hanford detector, the Livingston detector, and the Virgo detector, respectively. As pointed out later, three events fail the consistency check when comparing the npE posteriors from different detectors, suggesting certain noise artifacts that may be significantly affecting the inference process. These events are removed from the results presented in the main text. In addition, the strain data for GW190412 and GW250114_082203 is known to have significant contributions from higher harmonics, and so we take special care when modeling these signals.
By default, we choose IMRPhenomPv2 as the base GR waveform $\tilde{h}_{\rm GR}$. However, in the special cases of GW190412 and GW250114_082203, we choose IMRPhenomXPHM with an additional $(3,3)$ mode added on top of the dominant $(2,2)$ mode, as mentioned in the caption of Table 1. Both GR waveforms are parametrized by
@@
Another concern is that the naïve Gaussian KDE can extend outside the individual-event npE prior range, and, hence, affect the numerical result. To address this problem, we have run a diagnostic analysis where both the Gaussian KDE and the population model are truncated at $\vec{\zeta}=1$. We have checked that the posteriors of our hierarchical inference does not change under the truncation, implying that the naïve Gaussian KDE applied here is numerically robust.
@@
Figure 3: Posterior 90% credible contours from individual-event npE analysis assuming a uniform prior over $\zeta_{1}$ and $\zeta_{2}$ within the unit circle. Each panel presents the analysis result of an event in Table 1, as suggested by the panel title, and apart from the posterior generated using data of the entire network (blue), there are also posteriors generated using data of individual detectors including the Hanford detector (orange), the Livingston detector (green), and the Virgo detector (red). See Table 1 for the list of detectors used in each event. The ticks and labels of the panel axes have been omitted for simplicity. For all panels, the $x$-axis represents $\zeta_{1}\in[-1,1]$ and the $y$-axis represents $\zeta_{2}\in[-1,1]$. Each panel is also gridded by gray dotted lines, which are the same PN lines shown in Fig. 2 in the main text. The black “++” at the center of each panel marks GR.
@@
For the rest of the events, the network posteriors are mostly consistent with GR by enclosing $\vec{\zeta}=0$ inside the black 90% credible contours. Exceptions are GW191129_134029 and GW190924_021846, but their network 90% credible contours are not far from the GR point. There are also a few critical events such as GW150914, for which GR appears to be sitting on the 90% posterior contour. However, these cases should be seen as affected by instrument noise that happened to behave like non-GR deviations. This argument has been further strengthened by the hierarchical npE test result, from which we have presented a $Q_{\rm GR}=34\%<90\%$ in the main text. Here, the consistency with GR at the population level can be interpreted as a result from averaging out the random noise effects of individual events.
@@
Figure 4: Hierarchical posterior of the population hyperparameters. The blue contours enclose the 50% and 90% credible regions, respectively. The black dot and lines mark GR ($\mu=0=\sigma$).